OpenStudy (anonymous):

solve the equation in the interval [0,2pi)

OpenStudy (anonymous):

\[4\sin \theta-3=0\]

OpenStudy (anonymous):

Add 3 to the right side, and then divide by 4, so sin theta =3/4

OpenStudy (anonymous):

Rearrange the equation to get 4 sinX = 3

OpenStudy (anonymous):

Then rearrange to get sin X = 3/4

OpenStudy (anonymous):

i got that far

OpenStudy (anonymous):

Then take the arcsine of (3/4)

OpenStudy (anonymous):

If you use a calculator use (sin)^-1 to evaluate x in the range 0 to pi/2

OpenStudy (anonymous):

There should be 2 answers for the interval of 2pi

OpenStudy (anonymous):

Use the answer just found and subtract it from pi to get the second answer.

OpenStudy (anonymous):

ok i got thanks yall

OpenStudy (anonymous):

medal ?? ha ha

OpenStudy (anonymous):

@1on1 : You are welcome.

OpenStudy (anonymous):

@1on1 lol gotcha